OpenStudy (anonymous):

What is the solution of x + y = 2 and 2x + 5y = 16? x = 1, y = 4 x = 4, y = -2 x = -2, y = -2 x = -2, y = 4 Done

OpenStudy (anonymous):

let me show you

OpenStudy (anonymous):

x+y=2 2x+5y=16 x=2-y 2(2-y)+5y=16 4-2y+5y=16 3y=12 y=4 x+(4)=2 x=-2 (-2,4)

OpenStudy (jdoe0001):

same as show by Luigi0210 at http://openstudy.com/updates/52c1d35ee4b0fef2cb9f0ec1